OpenStudy (anonymous):

Would someone please help me? The volume in cubic feet of a workshop’s storage chest can be expressed as the product of its three dimensions: . The depth is x + 1. a. Find linear expressions with integer coefficients for the other dimensions. b. If the depth of the chest is 6 feet, what are the other dimensions?Answer

OpenStudy (anonymous):

do you have a picture of the volume? something appears to be missing

OpenStudy (anonymous):

No there is no picture The problem is v(x)=x^3-5x^2-2x+24. The depth is x+2.

OpenStudy (anonymous):

so really we want v(x) as a product of three linear factors

OpenStudy (anonymous):

oh i though there must be something else !

OpenStudy (anonymous):

I would assume correct. The word problem aga in v(x)=x^3-5x^2-2x+24. The depth is x +2. I need to find the linear expressions with integer coefficients for the other dimensions and if the depth of the chest is 75 feet what are the other dimensions as well? I am really lost here

OpenStudy (anonymous):

they are asking you to factor \[x^3-5x^2-2x+24=(x+2)\times \text{something}\]

OpenStudy (anonymous):

\[v(x)=x^3-5x^2-2x+24 = (x+2)(ax^2 + bx + c)\]

OpenStudy (anonymous):

eigen got this, i gotta run but that is the idea. divide to find the other factor

OpenStudy (anonymous):

Please show the work-I am so confused

OpenStudy (anonymous):

Are you still there eigenschmeigen?

OpenStudy (anonymous):

there are a few different approaches, but i find this the easiest to explain well we have a cubic v(x), and we want to express it as a product of three factors. a cubic can be written as the product of a linear factor and a quadratic, and that's a nice place to start: we'll use a,b,c to represent the coefficients of this quadratic. we are going to find a,b and c \[v(x)=x^3-5x^2-2x+24 = (x+2)(ax^2 + bx + c)\] expanding out on the right we have: \[(x+2)(ax^2 + bx + c) = ax^3 + bx^2 + cx + 2ax^2 + 2bx + 2c\] collect together the terms and put it back into the previous equation \[x^3-5x^2-2x+24 = ax^3 + (b + 2a)x^2 + (c+2b)x + 2c\] now because the has to work for all x we can compare coefficients: \[x^3 = ax^3 \text{ , } -5x^2 = (b+2a)x^2 \text{ , } -2x = (c+2b) \text{ , } 24 = 2c\] that first equation there looks pretty straight forward: \[x^3 = ax^3 \text{ , so: } a = 1 \] now the last equation: \[24 = 2c \text{ , so: } c = 12\] now lets look at the second equation there: \[-5x^2 = (b+2a)x^2\] \[-5 = b+2a\] since a = 1 \[-5 = b + 2\] \[b = -7\]

OpenStudy (anonymous):

that looks pretty long winded, once you have done many you need less and less of the method and it becomes quicker so we have: \[v(x)=x^3-5x^2-2x+24 = (x+2)(ax^2 + bx + c) = (x+2)(x^2 -7x +12)\]

OpenStudy (anonymous):

you can check your answer by expanding it back out. now we want to factorise the quadratic into two linear factors. are you ok doing that? \[v(x)= (x+2)(x^2 -7x +12)\]

OpenStudy (anonymous):

we get: v(x)= (x+2)(x -3)(x-4)

OpenStudy (anonymous):

Ok I am a little confused

OpenStudy (anonymous):

which bit is confusing for you? i'll try and explain more

OpenStudy (anonymous):

what would the linear expression with the integer coefficients be for the other dimensions? is that the b=-7?

OpenStudy (anonymous):

before we break it down into linear expressions we need to get the quadratic

OpenStudy (anonymous):

then we factorise the quadratic to get our other linear expressions

OpenStudy (anonymous):

I am reviewing your notes - v(x)= (x+2)(x -3)(x-4) is the linear espressions with coefficients for other dimensions?

OpenStudy (anonymous):

so depth is (x+2) length and width are (x-3) and (x-4)